Circular No. 8174

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2003 O3 (LINEAR) An apparently asteroidal object reported by LINEAR (discovery ob-

servation below), and posted on the NEO Confirmation Page, has been found to be apparently cometary on CCD images taken by P. Birtwhistle (Great Shefford, U.K., 0.30-m reflector; very faint tail ∼ 10′′ long in p.a. ≈ 270^{o}–280^{o} on July 31.10 and Aug. 2.08 UT; mag 18.1 and coma diameter ∼ 5′′ on Aug. 2.08), by J. Tichá and M. Tichý (Kleť, 1.06-m KLENOT telescope; diffuse with a wide tail in p.a. 260^{o} on Aug. 3.01), and by J. McGaha (near Tucson, AZ; possible tail spike 5′′ long in p.a. 300^{o} on Aug. 3.38 with a 0.30-m reflector; possible fan-shaped tail 5′′ long in p.a. 260^{o} on Aug. 5.33 with a 0.62-m reflector).

2003 UT  α_{2000}  δ_{2000}  Mag.
July 30.39211  1^{h}20^{m}09.^{s}26  +3^{o}40′44′′.0  18.9

The available astrometry, the following orbital elements (which indicate that the comet passed 0.3 AU from Jupiter in Nov. 1979), and an ephemeris appear on MPEC 2003-P25.

T = 2003 Aug. 13.955 TT  ω = 0.^{o}662
e = 0.63956  Ω = 342.106 }2000.0
q = 1.25725 AU  i = 8.583
a = 3.48812 AU n^{o} = 0.151292 P = 6.51 years

COMET P/2003 O2 (LINEAR) Additional astrometry, the following elliptical orbital elements, and an

ephemeris appear on MPEC 2003-P24.

T = 2003 Sept. 5.203 TT  ω = 32.^{o}338
e = 0.64895  Ω = 344.685 }2000.0
q = 1.50663 AU  i = 14.666
a = 4.29174 AU n^{o} = 0.110855 P = 8.89 years

COMET C/2002 O7 (LINEAR) Total visual magnitude estimates: June 18.96 UT, 12.5 (J. J. Gonzalez,

Leon, Spain, 0.20-m reflector); 21.92, 12.2 (J. Carvajal, Avila, Spain, 0.45m reflector); July 17.41, 11.4 (M. Mattiazzo, Wallaroo, South Australia, 0.28-m reflector); 27.40, 10.9 (Mattiazzo).

2003 August 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green